1.square root of 20 + square root of 80
2. square root of 12 - 2 square root of 3
3. sqare root of 28 - square root of 62
4. 3/square root of 6
5. square root of 24+ square root of 54+ 8 square root of6
6. square root of 32 + 2 square root of 2

Answer by checkley71(8403)

I'LL DO THE FIRST 3 TO SHOW YOU HOW TO DO THE REST. OK?
1) SQRT20+SQRT80=SQRT4*5+SQRT16*5=2SQRT5+4SQRT5=6SQRT5 ANSWER.
2) SQRT12-2SQRT3=SQRT4*3-2SQRT3=2SQRT3-2SQRT3=0 ANSWER.
3) SQRT28-SQRT62=SQRT4*7SQRT62=2SQRT7-SQRT62 ANSWER.
GOOD LUCK.